HUBBLE SPACE TELESCOPE DAILY REPORT #1750 PERIOD COVERED: 0000Z (UTC) 10/30/96 - 0000Z (UTC) 10/31/96 Daily Status Report as of 305/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 6541 (Interstellar Cadmium: Probing Galactic = Chemical Evolution) The High Resolution Spectrograph (Side Two) was used to make = spectrographic observations of the O/B class star HD24398 in order to = probe interstellar Cadmium and study the galactic chemical evolution. As = reported in yesterday's report and HSTAR 5769, two acquisitions and two = reacquisitions scheduled during this proposal failed due to a large saved = null bias error. Twelve observations were affected. No further problems = were reported. 1.2 Completed Two Sets of WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The acquisition and reacquisition problems mentioned in 1.1 also affected = the two observations in the first iteration of this proposal. Otherwise, = he observations were completed as planned, and no further problems were = reported. 1.3 Completed FOS/RD 6916 (FOS/Cycle 6:Photometric Calibration) The Faint Object Spectrograph (Red Side) was used to make = ultraviolet photometric calibration images of standard star BD+28D4211 = using the 4.3 arcsecond aperture. The loss of lock described in HSTAR = 5768, mentioned in yesterday's report, caused an FOS timeout and no = target acquisition occurred. Later in the proposal, another loss of lock = and another FOS timeout occurred. These latter events were recorded in = HSTARs 5771 and 5772. No further problems were reported. 1.4 Completed WF/PC-2 7203 (Cosmology with the Medium Deep Survey - = Short Orbit Cases) The WF/PC-2 was used to take short (about 20 minute) exposures, to = be taken in parallel with the prime science observations, using the I = filter. These exposures are part of the medium-deep survey, which = emphasizes exposures that will be useful in obtaining cosmological = parameters. The observations were completed as planned, and no problems = were reported. 1.5 Completed GHRS/1 6446 (The High Chromospheres of the late A Stars, = Altar and Alpha Cephei) The High Resolution Spectrograph (Side One) was used to make the = observations of the Si IV lines in the high chromosphere of the late, A = class, star HD203280. These data will be used to construct a model of = the chromosphere structure of this star. The observations were completed = as scheduled, and no anomalies were reported. At this point, the safemode event described in 2.3 occurred, and the following scheduled proposals did not execute: 1. FOS/RD 6577 (An Expanded Survey for Study of the Lyman- Alpha Line and QSO Mg II Absorption) 2. Two Sets of WF/PC-2 6253 (Pure Parallel Archive Program) 3. Two Sets of WF/PC-2 6909 (WF/PC-2 Cycle 6 Calibration: Earth Flats) 4. WF/PC-2 6344 (Low - Mass Binaries in the Hyades - Completing the Survey) 5. FOS/BL 6470 (Decoding the Inner Dsik About Beta-Pictoris) 6. COSTAR/U 6895 (FOC Cycle 6 PSF Monitoring) 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 1 Successful : 1 Scheduled Reacquisitions : 1 Successful : 1 2.2 FHST Updates: Scheduled : 2 Successful : 2 2.3 Operations Notes: On Wednesday, October 30, 1996, HST entered +V3 Software Sunpoint = Safemode at 304/181014Z. Safemode test 'First RWA Momentum Limit' on RWA = 1 failed. initiating the safemode event. AN format telemetry data was = restored at 304/181620Z. All spacecraft subsystems are healthy and = communications are nominal. Analysis of data is continuing in = preparation for an update briefing to HST Project Management to be held = today at 10:00 a.m. in Building 3, Room S166 when recovery procedures = will be discussed and finalized. Currently a Health and Safety Load is being prepared for uplink = later today. Others may follow. Normal science operations will resume = as determined by this morning's briefing. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH